-- Main.jill - 16 Jan 2007

-- Main.gredel - 18 Jan 2007

cirrus, high clouds moving in, low humidity 18%, how wind

pointer has been changed again to /Repository, let's see how this goes

00:17 25 x bias 00:38 dark 900 sec

00:55 opening ERROR same problem as last night, doors won't open left door opened manually

01:11 L95 Standard field U-band 1, 2, 10 sec almost completely overcast 5 x dither 10 sec U-band 5 x dither 20 sec U-band

01:30 L95 field B-band 5 x dither 20 sec

01:38 L95 field V-band 5 x dither 60 sec

01:49 L95 Standard B,V 1 and 2 sec each

01:53 image analysis

Computed Aberrations (nm) Defoc AstX AstY ComaX ComaY Sph (Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-767 457 -159 -715 -101 -312 372 344 165 -158 -115 -112 271 307 64 68 -80 -171 -6 -166 -374 223 -11 -52

02:03 Trilling 1st pass F1a --> image quality 1 arcsec in B F11

02:25 IA

02:30 Trilling F23 --> 1.18 arcsec F24 lbccontrol seems to hang, 'connecting to 10.144.0.21 -- waiting for 10.144.0.21

02:54 IQ degrades --> IA clouds are getting thicker

03:02 Trilling F56

lbccontrol hangs, timeout > 5 min --> reboot

03:23 image analysis, fails (no stars), scale exp time x 3 --> no pupils found

03:39 try Ryan 15deg_jan_epoch 1 --> sky at 16000 --> 1.9 arcsec IQ cannot do image analysis, might still be OK for Ryan

03:51 Ryan 10deg_jan_epoch 1 pretty cloudy, only a few stars visible

04:03 try IA again scale x 5 exposure time

Defoc AstX AstY ComaX ComaY Sph (Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
501 -664 379 2017 154 920 -1036 250 53 -1232 -63 -386 305 -430 94 -256 -162 -249

04:14 Trilling 1st field 2nd pass image quality 0.''9 F1a lbccontrol gui, relatively often 'hanging', message 'waiting for 10.144.0.21' displayed for several minutes F11 image elongation IA fails, scale x 5 exp time no pupils scale x 10 --> no pupils move telescope to clearer skies --> IA successful, will use that collimation for Trilling even if position differs

some discussions with Dave, image elongation (77 deg elevation) might be due to rotator tracking problems// or guiding? how many stars used for guiding?

[Note from John Hill the following day: The ~1 arcsec of image elongation seen in image 041716 appears to be astigmatism combined with varying focus across the field, and not any problem with the rotator or guider. I would say that IA dialed in some excess astigmatism prior to this exposure because of the ratty conditions.]

04:57 Trilling F11 F23 sky at 19000 F24 no stars F24 repeated --> getting better F56 better

05:56 preset to Ryan 15 deg, IA lbccontrol gui hanging again, waiting for something reboot after > 5 min waiting

information about network traffic would be most useful, not clear what is going on after reboot, TCS connection error --> lbckill lbcstart on cmu and restart csq

IA no pupils (but lots of clouds) -> scale exp time by 5; by 10 --> nothing trying...

06:44 succssful IA

Computed Aberrations (nm) Defoc AstX AstY ComaX ComaY Sph (Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
976 454 -635 1036 325 635 -614 193 -68 -577 -333 -204 589 -82 -100 -232 106 -55

--> 1.1 arcsec IQ

too late for 2nd epch for Ryan, too much time lost to clouds

06:53 Ryan 15deg_jan_epoch1 lbccontrol waiting again for something,

* 2 x 240 sec took 15 min to complete *

sky improving somewhat in NW, handful of stars visible

07:08 Trilling F297a 1st pass 2nd field -- looking through the clouds 07:23 F339 --> 15 min 07:31 F351a --> 8 min 07:40 F386a --> 9 min --> sky 16000 ADU F377a

preset to Ryan for IA waiting again for something - what is going on here??? IA --> fails; exposure time scale x 3, scale x 10 --> no pupils move to trilling again for collimation lbccontrol gui waiting for godot - sorry for 10.144.0.21 ping lbccontrol --> alive --> recovers again after 4 min waiting

Computed Aberrations (nm) Defoc AstX AstY ComaX ComaY Sph (Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
610 310 359 642 -198 701
  1. 321 204 889 -42 766 -980 19 -194 -355 173 -370 No GOOD pupils found... check input image!


d'accord, will try with last correction

08:33 Ryan 15deg_jan_epoch2 ryan 2 using correction from trilling field --> 1.3 arcsec IQ

08:47 Kochanek M81 V try old mirror settings first, if too bad, will do a IA where is M81? gone very slow again, waiting, connection, transferring data, waiting, connecting, transferring data from 10.144.0.21

--> looks like frequent connection failures to 10.144.0.21 suddenly going again
Computed Aberrations (nm) Defoc AstX AstY ComaX ComaY Sph (Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-503 -65 -73 -489 16 -283

09:03 Trilling F297 2nd pass --> no stars,

09:13 CLOSING, rainshower approaching rapidly 09:46 OPENING 'rain' (from radar) over - did not really rain, stayed all up in the atmosphere

09:51 Trilling 2nd pass F297a repeated F339 fails to do IA, no pupils, try exp. time scaling by 10 Computed Aberrations (nm) Defoc AstX AstY ComaX ComaY Sph (Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
231 73 -380 -81 247 -194

10:19 OB started on F351a skip B-band, only V hanging again after starting OB

* 10:24 rotator still in 'preset' * --> note in issue track filed

abort 10:33 OB restarted only sky -->

10:44 once again M81 ROTATOR IN PRESET AGAIN No error message in LogAnalyzer

abort again

10:50 just playing, try one of Peter's SN SN2006jc rotator went into preset again

10:55 CLOSING

Topic revision: r7 - 19 Jan 2007, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback